Solve for x.
Keep steps and answer in exact form.
(1/27)^x=9^(x-1)

Answer by Boreal(15235)

(1/27)^x=3^2(x-1)
3^(-3x)=3^2(x-1)
-3x=2x-2
5x=2
x=0.4